OpenStudy (anonymous):

Given sinθ = 5/9, where π/2 < θ < π, find the exact values of sin 2θ and cos 2θ.

OpenStudy (anonymous):

Im working it out again, but is the correct answer 10/9, -(4√14)/9

OpenStudy (campbell_st):

start by drawing a triangle in the 2nd quadrant  fin x, its negative... then you can find cos(@) and finally substitute the values into \[\sin(2\theta) = 2 \sin(\theta)\cos(\theta)\] and \[\cos(2\theta) = \cos^2(\theta) - \sin^2(\theta)\] hope it helps

OpenStudy (campbell_st):